A bond that matures in 15 years has a $1000 par value. The annual coupon interest rate is 8 percent and the market's required yield to maturity on a comparable-risk bond is 15 percent. What would be the value of this bond if it paid interest annually? What would be the value of this bond if it paid interest semiannually?
Answer #1

Details of Bond

Face Value $1000
Coupon Rate 8%
YTM 15%
Maturity 15years
Coupon Payment Annually 8%*1000 =$80
Coupon Payment Half Yearly 8%*100/2 = $40

Value of Bond if Coupon Payment is Annual

Value = 80(P/A,15%,15) +1000(P/F,15%,15)

from Compound interest table , value of P/A and P/F are taken

Value= 80*5.847 + 1000*0.122 = $589.76

Value of Bond if Coupon Payment is Semi Annual

In case of semiannual the interest rate is divided by 2 as in one year now due to semiannual calculation there will be 2 compounding periods and as a result time period gets doubled

Value = 40(P/A,15%/2,15*2) +1000(P/F,15%/2,15*2)

Value = 40(P/A,7.5%,30) +1000(P/F,7.5%,30)= 40*11.810 + 1000*0.144 = $616.4